What percent of 125 is 112.3?
Answer by ewatrrr(24785)

 
Hi
What percent of 125 is 112.3?
Let x represent the deciaml value: x*100 is the percent
x*125 = 112.3
x = 112.3/125
x = .8984
.8984*100 = 89.84%
CHECKING our Answer***
.8984*125 = 112.3